can someone help me with this problem, by completing each order of pair of the given equation
y =2x + 5: (8, ), (-1, ), ( , -1)

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)


Ordered pairs are called ordered because the values are given in a particular order. If the equation is in x and y, then the ordered pair is (x, y). So where you are given the first number in the ordered pair, you substitute that value for x in the equation and do the indicated arithmetic to determine the value for y that completes the ordered pair. Where you are given the second number in the ordered pair, substitute that number for y in the equation and do the indicated arithmetic to determine the value for x.


Answer by stanbon(75887)

can someone help me with this problem, by completing each order of pair of the given equation
y =2x + 5:
(8, ), : y = 2*8+5 = 21
(-1, ),: y = 2*-1+5 = 3
------------------------------
x = (y - 5)/2
( , -1): x = (-1 - 5)/2 = -3
================================
Cheers,
Stan H.